OpenStudy (anonymous):

What did I do wrong? Identify the vertex \[y=2x^2-4x-2\]\[=2(x^2-2x-1)\]\[=2(x-1)^2\] h=1 k=0 but according to my answers \[(h,k)\cancel{=}(1,0)\]

OpenStudy (anonymous):

@zepdrix

jimthompson5910 (jim_thompson5910):

x^2-2x-1 does NOT factor to (x-1)^2

jimthompson5910 (jim_thompson5910):

Notice this (x-1)^2 = (x-1)(x-1) (x-1)^2 = x(x-1)-1(x-1) (x-1)^2 = x^2 - x - x + 1 (x-1)^2 = x^2 - 2x + 1

jimthompson5910 (jim_thompson5910):

So if it was x^2 - 2x + 1, then it would factor to (x-1)^2

OpenStudy (anonymous):

Completing the square?

jimthompson5910 (jim_thompson5910):

yes you need to complete the square for x^2-2x-1

OpenStudy (anonymous):

I have a random question...I think I lost an excel document I worked very hard on. I stored it on my usb...Is that possible?

jimthompson5910 (jim_thompson5910):

its possible, and if you lost it you can search for it or you may be able to recover it if you deleted it on accident

OpenStudy (anonymous):

how?

jimthompson5910 (jim_thompson5910):

do a search for *.xls files

jimthompson5910 (jim_thompson5910):

that will list all the excel documents on your computer

OpenStudy (anonymous):

I have an outdated version of the stuff I worked on...It has a copy of 1/13/2013 I did so much more work on it since then...Oh well...Ok back to completing the square

jimthompson5910 (jim_thompson5910):

if you do that search, and that file exists somewhere, then you'll find it

OpenStudy (anonymous):

I worked on it on a computer at work, and the computer there give me a warning stating that it had protected information and wouldn't allow me to save it on my usb drive...I disregarded that message...do you think the computer would remove it from my usb?

OpenStudy (anonymous):

*gave me warning

jimthompson5910 (jim_thompson5910):

well if it said it couldn't save, then chances are that it never saved it in the first place

OpenStudy (anonymous):

oh I see...would it still be on that computer tomorrow?

jimthompson5910 (jim_thompson5910):

so it was probably all in memory (ie RAM) the whole time

jimthompson5910 (jim_thompson5910):

not if it only existed in RAM only

jimthompson5910 (jim_thompson5910):

although, there might be a chance it's on your work computer somewhere in temp file maybe

OpenStudy (anonymous):

yeah I hope so...I didn't shut the computer down.

jimthompson5910 (jim_thompson5910):

oh is the file/program still running?

OpenStudy (anonymous):

well I closed that (x-ed out of it) I'm such a noob when it comes to computer programs haha

jimthompson5910 (jim_thompson5910):

then chances are the data is gone because memory that is deallocated is given to another open program

OpenStudy (anonymous):

sigh.... <- sad koala.... how about a little completing the square exercise

jimthompson5910 (jim_thompson5910):

2x^2-4x-2 2(x^2-2x-1) 2(x^2-2x + 1 - 1 -1) 2((x^2-2x + 1) - 1 -1) 2((x-1)^2 - 2) 2(x-1)^2 - 4 Vertex: (1,-4)

OpenStudy (anonymous):

I recreated what I made at work. I saved it on my laptop this time!

jimthompson5910 (jim_thompson5910):

sounds good, hopefully it didn't take too long to recreate